# TextSplat: Text-Guided Semantic Fusion for Generalizable Gaussian Splatting

### v1

@@ Abstract.
Recent advancements in Generalizable Gaussian Splatting have enabled robust 3D reconstruction from sparse input views by utilizing feed-forward Gaussian Splatting models, achieving superior cross-scene generalization. However, while many methods focus on geometric consistency, they often neglect the potential of text-driven guidance to enhance semantic understanding, which is crucial for accurately reconstructing fine-grained details in complex scenes. To address this limitation, we propose TextSplat—the first text-driven Generalizable Gaussian Splatting framework. By employing a text-guided fusion of diverse semantic cues, our framework learns robust cross-modal feature representations that improve the alignment of geometric and semantic information, producing high-fidelity 3D reconstructions. Specifically, our framework employs three parallel modules to obtain complementary representations: the Diffusion Prior Depth Estimator for accurate depth information, the Semantic Aware Segmentation Network for detailed semantic information, and the Multi-View Interaction Network for refined cross‑view features. Then, in the Text-Guided Semantic Fusion Module, these representations are integrated via the text-guided and attention-based feature aggregation mechanism, resulting in enhanced 3D Gaussian parameters enriched with detailed semantic cues. Experimental results on various benchmark datasets demonstrate improved performance compared to existing methods across multiple evaluation metrics, validating the effectiveness of our framework. The code will be publicly available.
3D Reconstruction, Scene Understanding
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27-31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2025/06
@@ 1. Introduction
Recent advances in text‑driven vision–language models, such as CLIP (Radford et al., 2021) and Stable Diffusion (Rombach et al., 2022), have shown that integrating textual cues can markedly improve semantic coherence across tasks like image generation, object detection, and cross‑modal understanding. Motivated by these findings, we introduce TextSplat, the first text‑driven GGS framework that dynamically fuses multi‑source semantic representations via multi-level text embeddings to enhance geometry–semantic consistency in sparse‑view reconstruction. Specifically, our method first employs three parallel modules—Diffusion Prior Depth Estimator, the Semantic Aware Segmentation Network, and the Multi-View Interaction Network—to obtain accurate depth information, detailed semantic information, and refined cross-view features, respectively. Then, under sentence‑level embedding guidance, the Text-Guided Semantic Fusion Module dynamically integrates these representations via semantic aggregation to enrich multi‑source features and then applies content refinement to recover fine‑grained details lost during fusion. By integrating high-level semantic descriptions with geometric reasoning, our TextSplat generates robust cross-modal feature representations and achieves precise geometry–semantic alignment. This is particularly beneficial in sparse-view scenarios, where explicit semantic cues are essential for resolving ambiguities in complex scene structures. Experimental results on the RealEstate10K (Zhou et al., 2018) and ACID (Liu et al., 2021b) benchmark datasets demonstrate that TextSplat achieves state‑of‑the‑art performance, significantly improving reconstruction quality over existing 3DGS‑ and NeRF‑based methods.
@@ 2. RELATED WORK
Multi-View Stereo (MVS) (Furukawa et al., 2015; Wang et al., 2021) stands as a foundational methodology in computer vision for recovering three-dimensional scene geometries through the analysis of multiple two-dimensional image perspectives. NeRF (Neural Radiance Fields) (Mildenhall et al., 2021) is a neural rendering technique that marked a significant breakthrough in computer graphics by representing 3D scenes as continuous volumetric radiance fields, parameterized with neural networks to predict radiance and density at arbitrary spatial locations and viewing directions. However, this method is computationally intensive (Müller et al., 2022; Barron et al., 2022), often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. In contrast, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) emerges as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. It introduces an explicit scene representation using millions of learnable 3D Gaussians distributed in space, which preserves the strong fitting capability of continuous volumetric radiance fields for high-quality image synthesis. This innovative representation effectively circumvents the computational overhead inherent in NeRF-based methods, such as computationally expensive ray-marching operations and unnecessary calculations in empty spatial regions.
But real-world adoption of original NeRF and 3DGS pipelines faces critical bottlenecks due to their reliance on exhaustive multi-view capture. This challenge has spurred extensive research interest in sparse-view scene reconstruction and synthesis (Charatan et al., 2024; Chen et al., 2021, 2023; Gao et al., 2024; Niemeyer et al., 2022; Szymanowicz et al., 2024; Truong et al., 2023; Wu et al., 2024; Xu et al., 2024a; Zhang et al., 2024a), which seeks to generate high-fidelity 3D representations through limited viewpoint sampling. Current methodologies addressing sparse-view scenarios are systematically classified into two principal computational frameworks: scene-specific optimization paradigms and neural network-based feed-forward inference architectures. While the former necessitates time-consuming iterative optimization and exhibits limited generalization due to scene-specific parameter dependencies, the latter demonstrates superior computational efficiency and cross-scenario adaptability through its single-pass neural network pipelines and data-driven prior integration mechanisms.
@@ 3.1. Problem Definition
As illustrated in Fig. 2, we introduce TextSplat, a text-driven Generalizable Gaussian Splatting framework. Given $K$ sparse-view images =$\mathcal{I}=\{{\bm{I}}_{i}\}_{i=1}^{K}$ (${\bm{I}}_{i}\in\mathbb{R}^{H\times W\times 3}$) and their corresponding camera projection matrices $\mathcal{P}=\{{\bm{P}}_{i}\}_{i=1}^{K}$, our TextSplat predicts the comprehensive 3D Gaussian parameters $\mathcal{G}$. The camera projection matrix on the $i$-th view is: ${\bm{P}}_{i}={\bm{K}}_{i}[{\bm{R}}_{i}|\bm{t}_{i}]$, where ${\bm{K}}_{i}$, ${\bm{R}}_{i}$, and $\bm{t}_{i}$ are respectively the intrinsic, rotation and translation matrices. The 3D Gaussian parameters $\mathcal{G}$ include position $\mu$, opacity $\alpha$, covariance $\Sigma$, and color $c$. These parameters $\mathcal{G}$ support high-fidelity novel view synthesis by leveraging rasterization-based rendering of 3D primitives.
@@ 3.3. Semantic Aware Segmentation Network
The lack of viewpoints might hinder the holistic scene understanding and fine-grained perception. To handle this issue, we aim to incorporate the segmentation priors that can provide object-level semantic cues and holistic scene layout information. Following previous segmentation paradigms (Ronneberger et al., 2015; Cheng et al., 2022; Zhou et al., 2024; Li et al., 2024), we design the Semantic Aware Segmentation Network (SASN) with an encoder-decoder architecture, which effectively captures multi-scale semantic cues while preserving spatial-semantic consistency.
@@
Figure 3. Detailed Architecture of DPDE, SASN, and MVIN.
@@ 3.4. Multi-View Interaction Network
To fully exploit the complementary information across viewpoints and enhance the cross-view consistency for robust 3D reconstruction, we adopt a hybrid CNN-Transformer architecture (Xu et al., 2023, 2022) that models cross-view visual features through explicit inter-view interaction mechanisms. Specifically, we first employ a shallow ResNet-like CNN to extract per-view image features $\mathcal{CF}=\{CF_{i}\}_{i=1}^{K}$, followed by cross-view attention with self- and cross-attention layers to exchange information across different views as shown in Fig. 3 (c). Notably, the Transformer architecture specifically employs Swin Transformer’s local window attention (Liu et al., 2021a) to balance computational efficiency and global context modeling. This hybrid processing network ultimately generates cross-view semantic features =$\mathcal{MF}=\{MF_{i}\}_{i=1}^{K}$ that integrate both local texture patterns and global geometric consistency.
@@ 4.1. Experiment Setup
Model Initialization. The Multi-View Interaction Network and Cost Volume-based Decoder are initialized with pre-trained parameters from MVSplat (Chen et al., 2024b); the Diffusion Prior Depth Estimator inherits parameters from TADP (Kondapaneni et al., 2024); while the Semantic Aware Segmentation Network adopts weights from OMG-SEG (Li et al., 2024).
@@ 4.2. Comparison With State-of-The-Art
Table 4. Ablation Study of Core Components for TextSplat on RealEstate10K. Removing any key component leads to significant performance degradation.
@@
Table 5. Ablation Study of Training Stages on RealEstate10K, where TS denotes Training Stage. The progressive training strategy allows our model to fully realize its performance potential.

### v2

@@ Abstract.
Recent advancements in Generalizable Gaussian Splatting have enabled robust 3D reconstruction from sparse input views by utilizing feed-forward Gaussian Splatting models, achieving superior cross-scene generalization. However, while many methods focus on geometric consistency, they often neglect the potential of text-driven guidance to enhance semantic understanding, which is crucial for accurately reconstructing fine-grained details in complex scenes. To address this limitation, we propose TextSplat—the first text-driven Generalizable Gaussian Splatting framework. By employing a text-guided fusion of diverse semantic cues, our framework learns robust cross-modal feature representations that improve the alignment of geometric and semantic information, producing high-fidelity 3D reconstructions. Specifically, our framework employs three parallel modules to obtain complementary representations: the Diffusion Prior Depth Estimator for accurate depth information, the Semantic Aware Segmentation Network for detailed semantic information, and the Multi-View Interaction Network for refined cross-view features. Then, in the Text-Guided Semantic Fusion Module, these representations are integrated via the text-guided and attention-based feature aggregation mechanism, resulting in enhanced 3D Gaussian parameters enriched with detailed semantic cues. Experimental results on various benchmark datasets demonstrate improved performance compared to existing methods across multiple evaluation metrics, validating the effectiveness of our framework. The code will be publicly available.
@@ 1. Introduction
Recent advances in text‑driven vision–language models, such as CLIP (Radford et al., 2021) and Stable Diffusion (Rombach et al., 2022), have shown that integrating textual cues can markedly improve semantic coherence across tasks like image generation, object detection, and cross‑modal understanding. Motivated by these findings, we introduce TextSplat, the first text‑driven GGS framework that dynamically fuses multi‑source semantic representations via multi-level text embeddings to enhance geometry–semantic consistency in sparse‑view reconstruction as illustrated in Fig. 1. Specifically, our method first employs three parallel modules—Diffusion Prior Depth Estimator, the Semantic Aware Segmentation Network, and the Multi-View Interaction Network—to obtain accurate depth information, detailed semantic information, and refined cross-view features, respectively. Then, under sentence‑level embedding guidance, the Text-Guided Semantic Fusion Module dynamically integrates these representations via semantic aggregation to enrich multi‑source features and then applies content refinement to recover fine‑grained details lost during fusion. By integrating high-level semantic descriptions with geometric reasoning, our TextSplat generates robust cross-modal feature representations and achieves precise geometry–semantic alignment. This is particularly beneficial in sparse-view scenarios, where explicit semantic cues are essential for resolving ambiguities in complex scene structures. Experimental results on the RealEstate10K (Zhou et al., 2018) and ACID (Liu et al., 2021b) benchmark datasets demonstrate that TextSplat achieves state‑of‑the‑art performance, significantly improving reconstruction quality over existing 3DGS‑ and NeRF‑based methods.
@@ 2. RELATED WORK
Deep learning has been widely applied in various scenarios such as intelligent transportation (Li et al., 2021a, b; Yan et al., 2025), cross-modal learning (Liu et al., 2025; Zhao et al., 2025; Li et al., 2025), privacy protection (Li et al., 2023a, b), and artificial intelligence generated content (Li et al., 2024b; Liu et al., 2024). Within the domain of computer vision, Multi-View Stereo (MVS) (Furukawa et al., 2015; Wang et al., 2021) stands as a foundational methodology for recovering three-dimensional scene geometries through the analysis of multiple two-dimensional image perspectives. NeRF (Neural Radiance Fields) (Mildenhall et al., 2021) is a neural rendering technique that marked a significant breakthrough in computer graphics by representing 3D scenes as continuous volumetric radiance fields, parameterized with neural networks to predict radiance and density at arbitrary spatial locations and viewing directions. However, this method is computationally intensive (Müller et al., 2022; Barron et al., 2022), often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. In contrast, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) emerges as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. It introduces an explicit scene representation using millions of learnable 3D Gaussians distributed in space, which preserves the strong fitting capability of continuous volumetric radiance fields for high-quality image synthesis. This innovative representation effectively circumvents the computational overhead inherent in NeRF-based methods, such as computationally expensive ray-marching operations and unnecessary calculations in empty spatial regions.
But real-world adoption of original NeRF and 3DGS pipelines faces critical bottlenecks due to their reliance on exhaustive multi-view capture. This challenge has spurred extensive research interest in sparse-view scene reconstruction and synthesis (Charatan et al., 2024; Chen et al., 2021, 2023; Gao et al., 2024; Niemeyer et al., 2022; Szymanowicz et al., 2024; Truong et al., 2023; Wu et al., 2024; Xu et al., 2024a; Zhang et al., 2024a; Lin et al., 2024), which seeks to generate high-fidelity 3D representations through limited viewpoint sampling. Current methodologies addressing sparse-view scenarios are systematically classified into two principal computational frameworks: scene-specific optimization paradigms and neural network-based feed-forward inference architectures. While the former necessitates time-consuming iterative optimization and exhibits limited generalization due to scene-specific parameter dependencies, the latter demonstrates superior computational efficiency and cross-scenario adaptability through its single-pass neural network pipelines and data-driven prior integration mechanisms.
@@ 3.1. Problem Definition
As illustrated in Fig. 2, we introduce TextSplat, a text-driven Generalizable Gaussian Splatting framework. Given $K$ sparse-view images $\mathcal{I}=\{{\bm{I}}_{i}\}_{i=1}^{K}$ (${\bm{I}}_{i}\in\mathbb{R}^{H\times W\times 3}$) and their corresponding camera projection matrices $\mathcal{P}=\{{\bm{P}}_{i}\}_{i=1}^{K}$, our TextSplat predicts the comprehensive 3D Gaussian parameters $\mathcal{G}$. The camera projection matrix on the $i$-th view is: ${\bm{P}}_{i}={\bm{K}}_{i}[{\bm{R}}_{i}|\bm{t}_{i}]$, where ${\bm{K}}_{i}$, ${\bm{R}}_{i}$, and $\bm{t}_{i}$ are respectively the intrinsic, rotation and translation matrices. The 3D Gaussian parameters $\mathcal{G}$ include position $\mu$, opacity $\alpha$, covariance $\Sigma$, and color $c$. These parameters $\mathcal{G}$ support high-fidelity novel view synthesis by leveraging rasterization-based rendering of 3D primitives.
@@ 3.3. Semantic Aware Segmentation Network
Figure 3. Detailed Architecture of DPDE, SASN, and MVIN.
The lack of viewpoints might hinder the holistic scene understanding and fine-grained perception. To handle this issue, we aim to incorporate the segmentation priors that can provide object-level semantic cues and holistic scene layout information. Following previous segmentation paradigms (Ronneberger et al., 2015; Cheng et al., 2022; Zhou et al., 2024; Li et al., 2024a), we design the Semantic Aware Segmentation Network (SASN) with an encoder-decoder architecture, which effectively captures multi-scale semantic cues while preserving spatial-semantic consistency.
@@ 3.4. Multi-View Interaction Network
To fully exploit the complementary information across viewpoints and enhance the cross-view consistency for robust 3D reconstruction, we adopt a hybrid CNN-Transformer architecture (Xu et al., 2023, 2022) that models cross-view visual features through explicit inter-view interaction mechanisms. Specifically, we first employ a shallow ResNet-like CNN to extract per-view image features $\mathcal{CF}=\{CF_{i}\}_{i=1}^{K}$, followed by cross-view attention with self- and cross-attention layers to exchange information across different views as shown in Fig. 3 (c). Notably, the Transformer architecture specifically employs Swin Transformer’s local window attention (Liu et al., 2021a) to balance computational efficiency and global context modeling. This hybrid processing network ultimately generates cross-view semantic features $\mathcal{MF}=\{MF_{i}\}_{i=1}^{K}$ that integrate both local texture patterns and global geometric consistency.
@@ 4.1. Experiment Setup
Model Initialization. The Multi-View Interaction Network and Cost Volume-based Decoder are initialized with pre-trained parameters from MVSplat (Chen et al., 2024b); the Diffusion Prior Depth Estimator inherits parameters from TADP (Kondapaneni et al., 2024); while the Semantic Aware Segmentation Network adopts weights from OMG-SEG (Li et al., 2024a).
@@ 4.2. Comparison With State-of-The-Art
Table 4. Ablation Study of Core Components for TextSplat on RealEstate10K.
@@
4.3. Ablation Analysis
Table 5. Ablation Study of Training Stages on RealEstate10K, where TS denotes Training Stage.
@@ 5. Conclusion
Appendix A Qualitative Comparison of 3D Reconstruction Performance
As shown in Fig. 7, we have conducted qualitative visualization for depth estimation and 3D reconstruction. Specifically, within the stairwell region, our TextSplat approach achieves significantly more accurate depth estimation than the MVSplat. This superior performance translates to enhanced preservation of fine structural details within the reconstructed 3D Gaussian representations.
Figure 7. Qualitative Comparison of Depth Estimation 3D Reconstruction Performance.
Appendix B Feature Visualization
As shown in Fig. 8, we have conducted feature visualizations of the DPDE outputs, confirming that diffusion-based depth representations capture rich geometric patterns. These visualizations demonstrate that our method consistently generates high-fidelity depth cues across diverse scenes.
Figure 8. Feature Visualizations of the DPDE.
Appendix C Limitation
Despite the significant improvements achieved by our TextSplat framework in geometry-semantic consistency and reconstruction quality, certain limitations remain that warrant further investigation. As shown in Fig. 9, while our model effectively enhances reconstruction performance through its Text-Guided Semantic Fusion Module, challenges persist in accurately reconstructing scenes containing transparent materials such as glass doors and windows, as well as objects behind them. This limitation highlights the need for more sophisticated modeling of transparent materials and their interactions with surrounding objects. Future work will explore advanced material-aware priors and specialized loss functions to address these issues and further enhance scene fidelity.
Figure 9. Failure Cases on Transparent Materials.